Assume Mudvayne, Inc., has a 9.7 percent ROA and a 39 percent payout ratio.
Required:

What is the internal growth rate? (Do not include the percent sign (%). Round your answer to 2 decimal places (e.g., 32.16). )

  Internal growth rate %
Answer #1

Retention ratio=1-payout ratio

=(1-0.39)=0.61

Internal growth rate=(ROA*Retention ratio)/[1-(ROA*Retention ratio)]

=(0.097*0.61)/[1-(0.097*0.61)]

=6.29(Approx).